# HiReview: Hierarchical Taxonomy-Driven Automatic Literature Review Generation

### v1

@@
Abstract
In this work, we present HiReview, a novel framework for hierarchical taxonomy-driven automatic literature review generation. With the exponential growth of academic documents, manual literature reviews have become increasingly labor-intensive and time-consuming, while traditional summarization models struggle to generate comprehensive document reviews effectively. Large language models (LLMs), with their powerful text processing capabilities, offer a potential solution; however, research on incorporating LLMs for automatic document generation remains limited. To address key challenges in large-scale automatic literature review generation (LRG), we propose a two-stage taxonomy-then-generation approach that combines graph-based hierarchical clustering with retrieval-augmented LLMs. First, we retrieve the most relevant sub-community within the citation network, then generate a hierarchical taxonomy tree by clustering papers based on both textual content and citation relationships. In the second stage, an LLM generates coherent and contextually accurate summaries for clusters or topics at each hierarchical level, ensuring comprehensive coverage and logical organization of the literature. Extensive experiments demonstrate that HiReview significantly outperforms state-of-the-art methods, achieving superior hierarchical organization, content relevance, and factual accuracy in automatic literature review generation tasks.
1 Introduction
Figure 1: An illustration of automate literature review generation system.
Literature reviews play a crucial role in synthesizing knowledge from large bodies of work, helping to organize and summarize relevant research. However, manual literature reviews are labor-intensive and time-consuming, especially when addressing fields with complex, hierarchical topics. Consequently, there is growing interest in automating literature review generation (LRG) to reduce this burden. As shown in Fig. 1, an automated LRG system should accurately retrieve relevant papers, identify relationships between them, organize the papers and generate reliable, factually accurate content for the literature review.
Traditional multi-document summarization models (Erkan & Radev, 2004, Gunaratna et al., 2015, Izacard & Grave, 2020, Abdel-Salam & Rafea, 2022, Kasanishi et al., 2023) struggle to handle the extensive input lengths characteristic of scientific reviews, often resulting in shallow summaries that lack the depth required for comprehensive literature reviews. Recently, large language models (LLMs) have been applied to this task due to their ability to manage long-range dependencies and provide richer contextual understanding. However, directly employing LLMs in academic writing can lead to issues such as hallucinations—where the generated content is not grounded in factual data—and may fail to accurately reflect state-of-the-art research (Huang et al., 2023, Tonmoy et al., 2024). To mitigate these issues, some LLM-based approaches have integrated retrieval-augmented generation (RAG) to enhance content quality (Gao et al., 2023b, Fan et al., 2024). In LRG, tools like Paper Digest [^1] https://www.paperdigest.org/review/ can generate brief summaries but lack the depth and comprehensiveness required for detailed literature reviews. AutoSurvey (Wang et al., 2024) addresses the scalability challenges of traditional review methods by employing a specialized LLM in a multi-stage process that includes retrieval, chapter drafting, and refinement. Similarly, Wu et al. introduced an LLM-based method for automated review generation, consisting of four key components: literature retrieval, outline formulation, knowledge extraction, and review composition. However, these LLM-based approaches overlook critical prior knowledge, such as citation relationships, which are essential for retrieving relevant papers and producing coherent, accurate literature reviews. Additionally, they fail to organize papers into meaningful structures before the generation phase, limiting the system’s ability to fully comprehend the relationships between papers. These models rely too heavily on LLMs, with the generation process being primarily driven by prompts, resulting in less robust and often inconsistent outcomes.
To overcome the limitations of current LLM-based methods, we propose Hierarchical Taxonomy-Driven Automatic Literature Review Generation (HiReview), which fully leverages prior knowledge from citation networks, such as the network topology formed by papers, rather than relying solely on text. To enable LLMs to better understand the underlying relationships between papers, we propose generating a hierarchical taxonomy of the papers before proceeding with content generation. We introduce an end-to-end framework that generates a taxonomy from a large citation network. Specifically, we explore the relationships between papers in a hierarchical manner, where papers are divided into clusters, and each cluster at every hierarchical level corresponds to a specific topic in a field. We then employ a hierarchical generation strategy to determine the central topic for each cluster, ensuring coherence across the hierarchy. By organizing papers into a clear, structured taxonomy, HiReview enables LLMs to better comprehend the relationships between papers, resulting in robust, high-quality, and citation-aware literature reviews.
@@ 1 Introduction
  • Dataset. We construct a large-scale dataset of literature review papers. Each paper is annotated with its hierarchical taxonomy and related citation network, which serves as the foundation for training and evaluating literature review generation models.
  • Framework. We introduce HiReview, a novel hierarchical taxonomy-driven generation framework for literature review generation. Papers are first organized into a meaningful hierarchical taxonomy, which is then leveraged to guide the generation phase.
  • Robustness. We demonstrate the robustness of HiReview across various experiments. Compared to baseline models, HiReview exhibits superior consistency and quality in literature review generation, including better coverage, structure, and relevance.
2 Related Work
Retrieval-Augmented Generation (RAG). RAG has gained increasing attention for its ability to mitigate the hallucination problem in LLMs, thereby enhancing the trustworthiness of generated content (Lewis et al., 2020, Gao et al., 2023b, Chen et al., 2024). By dynamically retrieving relevant information from a large corpus during the generation process, RAG models improve the accuracy and relevance of outputs in applications such as citation-aware generation (Gao et al., 2023a, Menick et al., 2022), text evaluation (Yue et al., 2023, Xie et al., 2024), and open-domain question answering (Zhu et al., 2021, Izacard & Grave, 2020, Karpukhin et al., 2020). This approach helps ground the model’s output in up-to-date and relevant information, reducing the risk of hallucination and improving the reliability of large-scale generative models. More recently, graph retrieval-augmented generation (Hu et al., 2024, Edge et al., 2024, Peng et al., 2024) has garnered widespread attention, with relationships between documents or chunks proving to effectively enhance the scalability and performance of RAG in graph-related tasks (He et al., 2024, Mavromatis & Karypis, 2024).
@@ 2 Related Work
Parameter-Efficient Fine-Tuning (PEFT). Full-parameter fine-tuning of domain-specific large-scale pre-trained models requires significant resources. In contrast, parameter-efficient fine-tuning (PEFT) focuses on enhancing model performance in specific domains without updating all model parameters (Ding et al., 2023, Han et al., 2024). Key techniques in PEFT include Adapters, Low-Rank Adaptation (LoRA) (Hu et al., 2021), and Prompt Tuning. Adapters are small neural networks inserted into the layers of a pre-trained model to modify its behavior for specific tasks (Houlsby et al., 2019, Pfeiffer et al., 2020). Instead of updating the entire model, only the adapter parameters are updated, making the process computationally efficient. Adapters have also been explored in applying PEFT to graph-based language models (Chai et al., 2023, Perozzi et al., 2024, Liu et al., 2024). LoRA introduces low-rank matrices into transformer layers, enabling fine-tuning of only a small subset of the model’s parameters while preserving overall performance. Prompt Tuning involves learning task-specific prompts that guide the model to perform various tasks without altering its underlying parameters (Lester et al., 2021, Li & Liang, 2021, Liu et al., 2023). The pre-trained model remains fixed, with only the input prompts being optimized for the specific task.
3 Preliminaries
The goal of this paper is to generate a structured, summarized review in response to a user’s query. The process consists of three main stages: retrieving relevant papers, clustering papers based on shared characteristics, and generating a synthetic content summary. Before formally presenting the problem, we first define its key components to establish a shared understanding.
Citation Graphs. A citation graph is a directed graph where the nodes represent academic papers, and the edges represent the citation relationship between them. Formally, it can be defined as $G=(V,E,\{T_{u}\}_{u\in V})$, where $V$ denotes the set of nodes (i.e., scientific papers), $E\subseteq V\times V$ denotes the set of edges (i.e., citation relations). Specifically, $t_{u}\in T_{u}$ denotes the title of paper $u$. This representation enables the extraction of both structural and semantic information from the academic literature, forming the backbone of subsequent paper retrieval and clustering tasks.
Retrieval-Augmented Generation. To enhance the quality and relevance of the generated summary, we utilize a retrieval-augmented generation approach. Given a corpus of external knowledge sources $\mathcal{D}=\{d_{i}\}_{i=1}^{N}$ and a query $Q$, this method aims to leverage an optimal subset $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ to enrich the response generated by a PLM. Formally, the generative process is captured by the following probability distribution for the output sequence $Y$:
$p_{\Theta}(Y|Q,\mathcal{D})=\prod\nolimits_{i=1}^{r}p_{\theta}(y_{i}|y_{<i},X_{Q},[X_{d}]_{d\in\mathcal{D}^{\prime}}),$ (1)
where $\Theta$ denotes the PLM’s parameters, $Y$ is the token sequence in the generated response, $y_{<i}$ indicates the prefix tokens, $X_{Q}$ represents the token sequence of the query $Q$, and $[X_{d}]_{d\in\mathcal{D}^{\prime}}$ denotes the concatenation of token sequences from the relevant external sources retrieved from $\mathcal{D}^{\prime}$. This formulation ensures that the generated content is not only contextually accurate but also grounded in the retrieved academic literature.
Hierarchical Graph Clustering. A hierarchical graph $G=(V,E)$ is a multi-level representation in which nodes and edges are organized across various levels of granularity, leading to progressively abstract representations. Let $G_{l}=(V_{l},E_{l})$ denote the graph at level $l$ in the hierarchy, where $V_{l}$ and $E_{l}$ denote the set of nodes and edges at $l$-th level. The goal of hierarchical graph clustering (Xing et al., 2021, Ying et al., 2018, Ren et al., 2024) is to generate clusters $C_{l}=f(G_{l})$ at each level $l$, using a clustering algorithm $f(\cdot)$. At each level $l$, the clusters $C_{l}=\{c_{l}^{i}\}_{i=1}^{|V_{l+1}|}$ are formed, where each cluster $c_{l}^{i}\subset V_{l}$ represents a subset of nodes. These clusters $C_{l}$ at level $l$ serve as hyper-nodes for the next level $l+1$, forming a new graph $G_{l+1}=(V_{l+1},E_{l+1})$, with $V_{l+1}=C_{l}$. The process is applied recursively, aggregating clusters at each level until a stopping criterion is met, i.e., no further meaningful clusters can be identified.
Literature Review Generation (LRG). Given the hierarchical graph clustering approach, we now present the formal problem definition of LRGS. The task is to generate a comprehensive summary $\mathcal{R}$ based on a citation graph $G(V,E,\{T_{u}\}_{u\in V})$ and a research topic $Q$ as a query. The goal is to output =$\mathcal{R}=\{R_{k}\}_{k=1}^{K}$, where each $R_{k}$ is represented as a sentence or paragraph that discusses the most relevant papers to the query. A high-quality literature review must exhibit a clear structure, often reflected through an underlying taxonomy that categorizes technical topics within the field. To achieve this, we propose a two-stage taxonomy-then-generation framework. In the first stage, a taxonomy tree is constructed, categorizing papers based on techniques and topics relevant to the query. In the second stage, detailed content is generated for each topic based on this taxonomy and the corresponding set of related papers. The quality of the review depends heavily on how well this taxonomy organizes the topics, enabling a coherent and comprehensive overview of the field.
This definition highlights several key challenges as follows:
Challenge 1: How to accurately retrieve the relevant structured information from vast amount of citation networks? In LRG, the retrieval process is challenging due to the large size and complexity of citation networks, which contain not only textual information from individual papers but also structural relationships (i.e., citations). Successful LRG requires jointly considering the textual and structural information to identifying the most relevant sub-network from a vast citation network.
Challenge 2: How can diversified documents be classified by identifying correlations between them, taking into account the synergy between their textual content and topological structure? The topology of the citation network plays a critical role in writing a literature review, as citations indicate the influence and relevance of papers within a research field. Papers with stronger connections in the citation network tend to be more closely related. Therefore, an effective LRG system must consider both the textual content of papers and their relationships within the network to accurately identify clusters of relevant literature and reveal underlying patterns.
Challenge 3: How to generate the taxonomy in a hierarchical manner? The process of generating a coherent taxonomy tree is inherently hierarchical, with each level of the taxonomy tree closely related to the others. LRG requires identifying relevant technical topics and organizing them in a hierarchical fashion, capturing both broad categories and fine-grained subtopics.
4 Method
Figure 2: HiReview (taxonomy-then-generation). (a) Given a literature review topic, the most relevant community in the citation network is retrieved. (b) Papers are hierarchically divided into different clusters. (c) The central topic of each cluster at every level is generated. (d) Finally, the content of the literature review is generated with the hierarchical taxonomy.
In this section, we present our taxonomy-then-generation framework (Fig. 2). First, we introduce a graph retrieval strategy to address Challenge 1, which aggregates neighbor relevance scores during retrieval (Section 4.1). Next, we propose an end-to-end hierarchical taxonomy tree generation model, consisting of hierarchical clustering and hierarchical generation. Specifically, we introduce a novel hierarchical graph clustering approach that considers relationships between nodes across different levels of the hierarchy to tackle Challenge 2 (Section 4.2.1), categorizing papers in a citation network hierarchically. Then, in Section 4.2.2, we propose a bottom-up iterative generation approach to determine the central topic of each cluster at every hierarchical level, ultimately forming a taxonomy to address Challenge 3. Finally, we leverage the hierarchical taxonomy to guide the literature review generation process, producing high-quality, citation-aware literature reviews.
4.1 Graph Context-Aware Retrieval
Given a citation graph $G=(V,E,\{T_{u}\}_{u\in V})$ and a literature review topic $Q$, our objective is to identify the most relevant papers for the literature review. Different from common retrieval task, which usually only consider the textual similarity between query and source documents individually, to achieve the goal of finding related research works, we need to simultaneously consider the textual level relevance and the citation pattern between papers. To address this issue, we propose our unique textual subgraph retrieval method. Specifically, for each paper $u\in V$, we compute a relevance score $R(u,Q)$ between its title $t_{u}\in T_{u}$ and the query $Q$ using BM25 (a widely used retriever based on sparse retrieval). Rather than relying solely on the individual attributes of each paper, we enhance relevance scoring by incorporating information from the paper’s neighbors in the citation graph as:
$\tilde{R}(u,Q)=R(u,Q)+\sum\nolimits_{v\in\mathcal{N}(u)}\alpha\cdot R(v,Q),$ (2)
where $\mathcal{N}(u)$ denotes the set of neighbors of $u$ in $G$, and $\alpha$ is a pre-defined weighting factor that controls the influence of neighboring papers. We empirically find that aggregating the relevance scores of neighbors leads to a significant improvement in retrieval accuracy. This approach is effective because the relevance of a paper’s neighbors offers valuable contextual information that enhances its overall relevance to the topic. Further details are provided in Appendix A.3. The top-$k$ nodes with the highest aggregated scores are selected to form the subset $V^{\prime}$. Subsequently, we construct the subgraph $G^{\prime}(V^{\prime},E^{\prime},\{T_{u}\}_{u\in V^{\prime}})$, where $V^{\prime}$ is the set of selected nodes, $E^{\prime}=\{(u,v)\in E\mid u,v\in V^{\prime}\}$. We retain only the edges between the selected top-$k$ papers, thereby preserving the citation relationships among the most relevant papers.
4.2 Hierarchical Taxonomy Generation
In a literature review, each topic at every level corresponds to a set of papers, which aligns with a cluster in the citation graph. Therefore, we approach hierarchical taxonomy generation through hierarchical graph clustering, followed by generating a central topic for each cluster.
4.2.1 Hierarchical Citation Graph Clustering
We define a clustering function $f(\cdot)$, which takes the graph $G_{l}$ and node features $X_{G_{l}}$ at level $l$ as input, and then produces clusters, i.e. $C_{l}=f(G_{l},X_{G_{l}})$, where each cluster corresponds to a subgraph $G^{\prime}_{l}\subset G_{l}$. We leverage a text encoder to convert the title of nodes to initialization text embedding as $X_{G^{\prime}}=\text{LM}(\{t_{u}\}_{u\in V^{\prime}})$ [^2] Specifically, SentenceBert (Reimers, 2019) is used to generate the text embeddings.. For each cluster at level $l$, treating it as a hyper-node in the $l+1$ level, we construct the graph at level $l+1$ and generate new node features through an aggregation function $g$, i.e., $G_{l+1}=g(f(G_{l},X_{G_{l}}))$. The process continues recursively until a stopping criterion is met, i.e., no further meaningful clusters can be formed.
Clustering Function $f(\cdot)$. At each level $l$, we first update the node embeddings using a $\text{GNN}_{\theta}$, aggregating the features of neighboring nodes, as incorporating the information from references in the citation network is crucial. Then, we predict the probability $\hat{p}_{uv}$ that two nodes belong to the same cluster using an MLPϕ followed by a softmax transformation, where the concatenation of node embeddings for nodes $u$ and $v$ serves as the input. We then calculate the node density, which measures the similarity-weighted proportion of same-cluster nodes in its neighborhood as,
$\hat{d}_{u}=\frac{1}{|\mathcal{N}(u)|}\sum\nolimits_{k\in\mathcal{N}(u)}\hat{p}_{uk}\cdot\frac{h_{u}\cdot h_{k}}{\|h_{u}\|\|h_{k}\|}.$ (3)

### v2

# Taxonomy Tree Generation from Citation Graph

@@
Abstract.
Constructing taxonomies from citation graphs is essential for organizing scientific knowledge, facilitating literature reviews, and identifying emerging research trends. However, manual taxonomy construction is labor-intensive, time-consuming, and prone to human biases, often overlooking pivotal but less-cited papers. In this paper, to enable automatic hierarchical taxonomy generation from citation graphs, we propose HiGTL (Hierarchical Graph Taxonomy Learning), a novel end-to-end framework guided by human-provided instructions or preferred topics. Specifically, we propose a hierarchical citation graph clustering method that recursively groups related papers based on both textual content and citation structure, ensuring semantically meaningful and structurally coherent clusters. Additionally, we develop a novel taxonomy node verbalization strategy that iteratively generates central concepts for each cluster, leveraging a pre-trained large language model (LLM) to maintain semantic consistency across hierarchical levels. To further enhance performance, we design a joint optimization framework that fine-tunes both the clustering and concept generation modules, aligning structural accuracy with the quality of generated taxonomies. Extensive experiments demonstrate that HiGTL effectively produces coherent, high-quality taxonomies.
Hierarchical taxonomy generation, citation graph, graph learning, retrieval-augmented generation, large language models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Graph algorithmsccs: Information systems Information extraction
1. Introduction
Figure 1. The citation graph and its taxonomy exhibit a clear hierarchical mapping.
Citation graphs represent the citation relationships within a collection of documents (e.g., research papers, patents, etc.), capturing the structure of scholarly communication and illustrating how ideas evolve and influence one another. A taxonomy of citation graphs organizes research topics hierarchically, grouping related papers into clusters that form broader categories (as shown in Figure 1). This tree-like structure highlights the connections between specific subtopics and overarching fields, offering an interpretable overview of the research landscape. Such taxonomies are typically substantial backbones of subsequent activities such as literature reviews (Shen et al., 2018a), meta-analysis and knowledge discovery (Hearst, 1992; Liu et al., 2012), and research trend identification (Ponzetto and Navigli, 2010), which are vital to the scientific community. However, manual taxonomy construction is labor-intensive, time-consuming, and prone to human bias and comprehensiveness, which hence struggle to keep pace with the rapid growth of publications. In natural language processing, there are techniques called existing taxonomy learning which focuses on deriving semantic relationships from linguistic patterns in texts (Hearst, 1992; Pantel and Ravichandran, 2004; Snow et al., 2006; Suchanek et al., 2007). They are designed for plain text corpora rather than text graphs and hence cannot handle patterns like connectivities and communities of documents, which are crucial for taxonomy generation from citation graphs (Liu et al., 2012; De Knijff et al., 2013; Dietz et al., 2012; Zhang et al., 2018). In this paper, we focus on a compelling problem:
Can we automatically generate taxonomies from citation graphs?
Generating taxonomy trees from citation graphs, though important, encompasses significant and unique major challenges to be addressed: (1) How do graph topology and texts jointly decide taxonomy? Citation graphs contain both 1) the citing relationships among papers with similar topics of interest and 2) the rich textual content of each paper. Effectively integrating these two modalities is crucial because relying solely on graph topology may overlook semantic nuances, while focusing only on text may ignore the technique evolution and contextual relevance indicated through citation links. (2) How are the taxonomy nodes verbalized? Verbalizing taxonomy nodes involves generating concise, accurate labels that capture the essence of each cluster while maintaining semantic coherence across hierarchical levels. This is challenging because clusters often span diverse or overlapping topics, requiring the model to balance specificity and generalization to produce informative and contextually appropriate labels. (3) How to automatically learn the whole taxonomy generator from data? Transforming from citation graphs to taxonomy is a nontrivial tasks that require a highly expressive model for both taxonomy tree construction and taxonomy node verbalization. How to train such a model from limited data is challenging in both the optimization and learning aspects.
To pursue our goal and address these challenges, we propose Citation Graph to Taxonomy Transformation (CGT), a novel framework that automatically extracts taxonomies from citation graphs. First, to enable meaningful hierarchical citation graph clustering, we propose a novel method that recursively decomposes the citation graph into multiple levels by grouping papers based on both textual content and citation link topology. At each level, nodes of lower-level topics are clustered into “supernodes” of higher-level topics, while feature aggregation maintains semantic coherence, facilitating the extraction of structured taxonomies. Second, to verbalize the nodes in taxonomy, we introduce a hierarchical concept abstraction strategy that iteratively generates central concepts for each cluster. The method synergizes cluster-level graph embeddings and paper-level information to produce consistent, semantically rich taxonomies aligned with user-preferred topics by prompting. Finally, to ensure clusters are structurally meaningful and semantically coherent, we jointly optimize all the parameters of our CGT through a customized training strategy consisting of pre-training and fine-tuning. To thoroughly evaluate the effectiveness of our method, we collected 518 citation graphs, each corresponding to a citation graph of a high-quality, human-written literature review in the computer science domain. Extensive experiments demonstrate that our model effectively generates high-quality taxonomies.
@@ 1. Introduction
  • We introduce a novel end-to-end framework that automatically generates taxonomies from citation graphs, guided by human-provided central topics to enhance relevance.
  • We propose a hierarchical citation graph clustering approach for citation graphs that integrates textual content and structural information, enabling the formation of semantically meaningful and structurally coherent clusters.
  • We develop a taxonomy node verbalization method that iteratively generates central concepts for clusters at each level, ensuring semantic consistency throughout the hierarchy.
  • We design a novel optimization framework that jointly fine-tunes hierarchical citation graph clustering and taxonomy node verbalizer, improving both structural accuracy and the coherence of the generated taxonomies.
2. Related Work
2.1. Taxonomy Learning
Taxonomy learning (Hearst, 1992; Pantel and Pennacchiotti, 2006; Suchanek et al., 2006) aims to construct hierarchical structures that capture semantic relationships within domain-specific corpora. Its objective is to identify and organize conceptual relationships while maximizing semantic coherence and relevance. Existing approaches can be categorized into four main paradigms: pattern-based methods (Suchanek et al., 2006; Ponzetto and Strube, 2011; Rios-Alvarado et al., 2013), clustering-based techniques (Liu et al., 2012; Dietz et al., 2012), statistical frameworks (Wang et al., 2009; Diederich and Balke, 2007), and graph-based approaches (Kozareva and Hovy, 2010; Velardi et al., 2013; Kang et al., 2015). While these methods have advanced our understanding of taxonomy construction, they predominantly focus on unstructured database (Huang et al., 2019; Qiu et al., 2018), limiting their applicability to more complex structured databases. Recent advances have extended taxonomy learning to graph-structured data, particularly knowledge graphs (Martel and Zouaq, 2021). However, these approaches are primarily designed for knowledge graphs where nodes represent atomic concepts or entities. This makes them insufficient for real-world graph databases like citation networks, where nodes contain rich textual content.
2.2. Hierarchical Graph Clustering
The primary goal of hierarchical graph clustering is to construct a hierarchy of clusters that effectively captures the underlying multi-scale structure of a graph. Flake et al. introduced a minimum s-t-cuts-based hierarchical graph clustering method, guaranteeing high-quality partitions. Doll et al. extended this approach by allowing arbitrary minimum s-t-cuts, enabling more flexible clustering hierarchies. While these methods provide strong theoretical guarantees, they struggle with scalability on large-scale graphs. To improve efficiency, density-based methods have been explored (Schlitter et al., 2014; Szilágyi et al., 2014), grouping nodes based on high local density, making them robust to variations in graph topology. Modularity-based methods have also been successful. Zeng and Yu proposed a parallel hierarchical graph clustering algorithm leveraging distributed memory architectures and a divide-and-conquer strategy for large-scale graphs. These methods remain widely used in unsupervised community detection (Bonald et al., 2018; Zeng and Yu, 2015). To improve scalability, minimum spanning tree (MST)-based clustering methods, such as Affinity Clustering (Elmahdy et al., 2020) and Hierarchical Spectral Clustering (Singh and Ganapathy, 2025), have been proposed to recursively partition graphs.
2.3. Parameter-Efficient Fine-Tuning (PEFT)
@@
Full fine-tuning of domain-specific large pre-trained models is resource-intensive, while parameter-efficient fine-tuning (PEFT) improves performance in specific domains by updating only a subset of parameters (Ding et al., 2023; Han et al., 2024). Key techniques in PEFT include Adapters, Low-Rank Adaptation (LoRA) (Hu et al., 2021), and Prompt Tuning. Adapters are small neural networks inserted into the layers of a pre-trained model to modify its behavior for specific tasks (Houlsby et al., 2019; Pfeiffer et al., 2020). Instead of updating the entire model, only the adapter parameters are updated, making the process computationally efficient. Adapters have also been explored in applying PEFT to graph-based language models (Chai et al., 2023; Perozzi et al., 2024; Liu et al., 2024). LoRA introduces low-rank matrices into transformer layers, enabling fine-tuning of only a small subset of the model’s parameters while preserving overall performance. Prompt Tuning involves learning task-specific prompts that guide the model to perform various tasks without altering its underlying parameters (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2023). The pre-trained model remains fixed, with only the input prompts being optimized for the specific task.
Figure 2. Overall framework of proposed hierarchical taxonomy generation.
3. Problem Formalization
Citation Graph. A citation graph is a directed graph defined as $\mathcal{G}=(V,E,\{D_{v}\}_{v\in V})$, where $V$ is the set of nodes (papers), $E\subseteq V\times V$ represents the directed edges (citation relationships), and $D_{v}=[w_{1},w_{2},\dots,w_{n}]$ denotes the tokenized textual content of paper $v$, where $w_{i}\in\mathcal{W}$, $\mathcal{W}$ represents the token vocabulary, and $n$ is the number of tokens in $D_{v}$.
Taxonomy Tree. A taxonomy tree, $\mathcal{T}=(U,R,\{D_{u}\}_{u\in U})$, is a rooted tree where $U$ is the set of nodes representing topics or subtopics, $R\subseteq U\times U$ denotes directed edges capturing hierarchical relationships, and $D_{u}=[w_{1},w_{2},\dots,w_{k}]$ represents the tokenized textual content associated with each topic $u$, where $w_{i}\in\mathcal{W}$ and $k$ is the number of tokens in $D_{u}$. The taxonomy tree outlines relationships among different topics in the citation graph, with each topic corresponding to a paper cluster.
Taxonomy Tree Generation from Citation Graph. The goal of taxonomy generation is to learn a mapping function $f$ that constructs an optimal taxonomy tree from the citation graph for a specific topic $q$, i.e., $f:\mathcal{G},q\rightarrow\mathcal{T^{*}}$. To achieve this goal, three challenges must be addressed:
  • Graph-to-Taxonomy Transformation. Citation graphs exhibit complex citation relationships, often allowing papers to belong to multiple topics, whereas taxonomy trees impose a strict hierarchical structure with a single-parent constraint.
  • Semantic Coherence in Taxonomy Generation. Citation graphs contain extensive textual data with redundancy across papers. Constructing a concise, coherent taxonomy requires effective topic partitioning while preserving semantic integrity.
  • The Learning of Taxonomy Generation. The optimization must ensure structural validity and semantic coherence by jointly refining tree structure and topic generation, aligning fine-grained clusters with higher-level generated concepts.
4. Methodology
4.1. Overview
Given a citation graph $\mathcal{G}=(V,E,\{D_{v}\}_{v\in V})$, taxonomy generation for $\mathcal{G}$ involves both structural organization and concept compression. Formally, let a hierarchy induction function $g:(V,E)\to(U,R)$ define the hierarchical organization of topics, and let a concept abstraction function $h:\{D_{v}\}_{v\in V}\to\{D_{u}\}_{u\in U}$ compress and summarize textual content at each taxonomy level. We argue that these mappings are inherently interdependent, forming an entangled transformation expressed as $f=h\circ g$ (Figure 2(a)), which decomposes the taxonomy generation process for a citation graph into two interlinked subproblems: Hierarchical Citation Graph Clustering and Taxonomy Node Verbalization. A well-designed hierarchy induction function $g$ establishes the relationship between the citation graph and its taxonomy. Subsequently, an effective concept abstraction function $h$ extracts key concepts and topics from the citation graph. To achieve this decomposition and address the challenges at the end of Section 3, we propose Hierarchical Graph Taxonomy Learning (HiGTL) for citation graph:
  • In Section 4.2, we propose a novel hierarchical graph clustering approach as the hierarchy induction function to address the Graph-to-Taxonomy Transformation challenge. We hierarchically categorize papers in a citation network, ensuring that each cluster at a given level corresponds to a topic in the taxonomy tree at the same level.
  • In Section 4.3, to achieve taxonomy node verbalization and address the challenge of Semantic Coherence in Taxonomy Generation, we propose an iterative graph-taxonomy generation approach—a bottom-up method that identifies the central topic of each cluster at every hierarchical level while ensuring conceptual alignment across levels.
  • In Section 4.4, we propose a novel two-phase optimization strategy for the Learning of Taxonomy Generation, enabling end-to-end joint training of graph clustering and taxonomy generation. The process begins with pretraining the hierarchical citation graph clustering module using discrete clustering labels, followed by fine-tuning both the clustering and generation modules with continuous labels representing topic content.
4.2. Hierarchical Citation Graph Clustering
Given a citation graph $\mathcal{G}$, let $\{\mathcal{G}^{(l)}=(V^{(l)},E^{(l)},X^{(l)})\}_{l=1}^{L}$ represent its hierarchical decomposition into a sequence of (hyper-)graphs across $L$ levels, where $V^{(l)}$ are the nodes, $E^{(l)}$ are the edges, and $X^{(l)}$ are the node representations at level $l$. The recursive updates are defined as:
(1) $(V^{(l+1)},E^{(l+1)})=\text{CLU}(V^{(l)},E^{(l)},X^{(l)}),$
(2) $X^{(l+1)}=\text{AGG}(\text{CLU}(V^{(l)},E^{(l)},X^{(l)})),$
where $\text{CLU}(\cdot)$ clusters nodes in $\mathcal{G}^{(l)}$ into hyper-nodes for $\mathcal{G}^{(l+1)}$ and forms edges between them (elaborated in Section 4.2.1), while $\text{AGG}(\cdot)$ computes features for hyper-nodes in $V^{(l+1)}$ (elaborated in Section 4.2.2). The sequence is initialized from the citation graph $\mathcal{G}^{(1)}=\mathcal{G}$, with node features $X^{(1)}=\text{LM}(\{D_{v}\}_{v\in V})$ generated using pre-trained dense language models.
The hierarchical citation graph clustering and aggregation operations collectively define the hierarchy induction function $g:(V,E)\to(U,R)$. This hierarchical citation graph clustering partitions the node set $V^{(l)}$ into clusters $\mathcal{C}^{(l)}$, with each hyper-node $u\in V^{(l+1)}$ corresponding to a unique topic $u\in U$. The edge set $R$ is directly induced from the evolving hierarchical aggregation across hierarchical levels.
4.2.1. Clustering Operator $\text{CLU}(\cdot)$.
The clustering function partitions nodes in $V^{(l)}$ into clusters based on structural and feature similarities, treating each cluster as a hyper-node in $V^{(l+1)}$. Edges $E^{(l+1)}$ are formed between hyper-nodes if any nodes in their corresponding clusters were connected in $E^{(l)}$:
$E^{(l+1)}=\{(c^{(l)}_{i},c^{(l)}_{j})\mid\exists(u,v)\in E^{(l)},\ u\in c^{(l)}_{i},\ v\in c^{(l)}_{j}\}.$
where $c^{(l)}_{i}$ and $c^{(l)}_{j}$ represent node clusters at level $l$. We next introduce our strategy for extracting clusters at each hierarchical level.
[…]
(3) $\hat{d}_{u}=\frac{1}{|\mathcal{N}(u)|}\sum_{k\in\mathcal{N}(u)}\hat{p}^{(l)}_{uk}\cdot\frac{h_{u}\cdot h_{k}}{\|h_{u}\|\|h_{k}\|}.$